OpenStudy (anonymous):

How many solutions does the system of equations have? X=6Y+6, 9X+3Y=18

OpenStudy (amoodarya):

put x=6y+6 in second equation to find y

OpenStudy (amoodarya):

9( 6Y+6)+3Y=18 54y+54+3y=18 /3 18y+18+y=6 19y=6-18 y=-12/19 so has one solution

OpenStudy (anonymous):

That's the same thing I got. Thanks